OpenStudy (yrivers36):

Please help solve with explanations: 1/2(x-1)+2x=3x+1

Directrix (directrix):

(1/2) (x-1) + 2 x = 3x + 1 Is this the problem?

Directrix (directrix):

Assuming that it is, then multiply through by 2. That gives (x - 1) + 4 x = 6x + 2 5x -1 = 6x +2 5x - 6x = 2 + 1 -x = 3 x = -3

OpenStudy (yrivers36):

Thank you so much

OpenStudy (yrivers36):

The answer should be -1/3